# Variational Mode Decomposition–Based Nonstationary Coherent Structure Analysis for Spatiotemporal Data

### v1

@@ Abstract
The conventional modal analysis techniques face difficulties in handling nonstationary phenomena, such as transient, nonperiodic, or intermittent phenomena. This paper presents a variational mode decomposition–based nonstationary coherent structure (VMD-NCS) analysis that enables the extraction and analysis of coherent structures in case of nonstationary phenomena from high-dimensional spatiotemporal data. The VMD-NCS analysis decomposes the input spatiotemporal data into intrinsic coherent structures (ICSs) that represent nonstationary spatiotemporal patterns and exhibit coherence in both the spatial and temporal directions. Furthermore, unlike many conventional modal analysis techniques, the proposed method accounts for the temporal changes in the spatial distribution with time. The performance of the VMD-NCS analysis was validated based on the transient growth phenomena in the flow around a cylinder. It was confirmed that the temporal changes in the spatial distribution, depicting the transient growth of vortex shedding where fluctuations arising in the far-wake region gradually approach the near-wake region, were represented as a single ICS. Further, in the analysis of the quasi-periodic flow field around a pitching airfoil, the temporal changes in the spatial distribution and the amplitude of vortex shedding behind the airfoil, influenced by the pitching motion of the airfoil, were captured as a single ICS. Additionally, the impact of two parameters, adjusting the number of ICSs ($K$) and the penalty factor related to the temporal coherence ($\alpha$), was investigated. The results revealed that $K$ has a significant impact on the VMD-NCS analysis results. In the case of a relatively high $K$, the VMD-NCS analysis tends to extract more periodic spatiotemporal patterns resembling the results of dynamic mode decomposition, whereas in the case of a small $K$, the analysis tends to extract more nonstationary spatiotemporal patterns.
@@
journal: preprint server
@@ 1 Introduction
Generally, many physical phenomena generate complex spatiotemporal structures owing to their nonlinearity. In recent years, techniques for extracting spatiotemporal patterns from high-dimensional time-series data obtained through computational fluid dynamics or experiments have been actively researched in the fluid dynamics field. These techniques aim to elucidate the physically significant fluid structures and their temporal behaviors within complex flow fields. Additionally, they are employed to model the flow field for prediction and control Taira2017 ; Rowley2017 ; Taira2020 ; Unnikrishnan2023 . The most common analytical methods are proper orthogonal decomposition (POD) Lumley1967 ; Berkooz1993 and dynamic mode decomposition (DMD) Schmid2010 . POD identifies the dominant spatiotemporal patterns by computing orthogonal bases that minimize the reconstruction error when reconstructing the original data. DMD represents the flow field’s temporal evolution from one time to the next using a linear operator, thereby extracting modes as eigenvalues and eigenvectors of the operator. Due to the linear-operator representation, each mode exhibits a linear growth in periodic oscillation amplitudes, simplifying the physical interpretation of the obtained modes. These techniques have found extensive applications in the extraction of spatiotemporal patterns from complex flow fields, such as jets Rowley2009 ; Jovanovic2014 , wakes Muld2012 ; Ohmichi2019 , and flows involving shockwaves Grilli2012 ; Ohmichi2018 . Moreover, substantial research efforts have been devoted to developing advanced methodologies, encompassing algorithms tailored for large and streaming datasets Hemati2014 ; Ohmichi2017c ; Amor2023 , noise-robust algorithms Dawson2016 ; Hemati2017 ; Nonomura2019 ; Sashidha2022 ; Ohmichi2022ttls , data-reconstruction methods from noisy Pastuhoff2013 ; Sugioka2019 ; Ohmichi2022 or incomplete Everson95 ; Nekkanti2023 data, and stability analysis methods Ranjan2020 ; Stahl2023 .
@@
However, the existing modal analysis techniques encounter difficulties in handling transient, nonperiodic, or intermittent phenomena (nonstationary phenomena). For instance, the DMD method represents each mode’s temporal behavior as a linear growth and periodic oscillation. Thus, it is unable to capture temporal variations in amplitude and frequency inherent in nonstationary phenomena. The POD method is frequently limited by the coexistence of various phenomena within a single mode, complicating its physical interpretation. Additionally, most existing modal-decomposition methods rely on input data representations of the following form Noack2016 :
$f(\bm{x},t)=\sum_{i=1}^{N}{c_{i}(t)\phi_{i}(\bm{x})}+\epsilon(\bm{x},t)$ (1)
@@
Here, the input data are represented as a superposition of $N$ space-dependent modes, $\phi_{i}$. The coefficients, $c_{i}$, are expansion coefficients that vary with time. $\epsilon$ denotes the expansion residual. Equation (1) indicates that the spatial distribution of mode $\phi_{i}$ remains unchanged with time. However, many nonstationary phenomena, such as the transient growth of a cylinder wake, involve spatial structures that vary with time. Therefore, to extract such phenomena as a single mode, a modal expansion using ‘flexible’ modes capable of changing spatial structures is required, as noted by Noack Noack2016 .
@@
In the field of signal processing, variational mode decomposition (VMD) Dragomiretskiy2014a and its multivariate counterpart, multivariate VMD (MVMD) Rehman2019 , have recently found applications as analytical tools for addressing nonstationary time-varying data. VMD tackles optimization problems by incorporating penalty terms related to the frequency bandwidth of modes, thereby extracting the coherent fluctuation patterns (modes) within the input signal. Unlike DMD, VMD does not assume single periodic variations and can represent nonlinear changes in amplitude. Each mode exhibits a distinct frequency spectrum with different center frequencies. Additionally, by extracting modes with narrow frequency bandwidths, it becomes feasible to isolate similar frequency variations as one coherent phenomenon, enhancing the ease of deriving physical interpretations. However, directly applying VMD algorithms to high-dimensional time-series data, such as fluid data, is challenging.
@@
Notably, only a few studies have attempted to develop modal analysis techniques for nonstationary fluid phenomena. The study by Liao et al. Liao2023 is the sole study focusing on an approach based on VMD. They proposed reduced-order VMD (RVMD) inspired by the Hilbert–Huang transform and VMD. However, RVMD is a modal-decomposition method that is represented as Equation (1), which cannot capture the temporal evolution of spatial distributions of nonstationary phenomena as a single mode. In the approach based on DMD, called multi-resolution DMD (mrDMD) Kutz2016 , different dynamic modes are extracted for each temporal domain of data based on the repetitive removal of low-frequency components and data segmentation. The conditional space–time POD method Schmidt2019 extracts dominant spatiotemporal patterns from a set of event realizations. The modes of the conditional space-time POD exhibit coherent structures in both spatial and temporal directions. Recently, Stahl et al. Stahl2023 reported that convective instability can be analyzed by applying DMD to conditional space-time POD modes.
@@ 2.1 Variational Mode Decomposition
where $K$ is the number of modes. VMD is a data-driven modal-decomposition method, and each mode, $u_{k}(t)$, is computed from the input signal, $y(t)$. Each mode is obtained by solving the following constrained variational optimization problem, aiming to minimize the bandwidth of the frequency distribution of each VMD mode and the reconstruction error.
@@
Here, $\left(\delta(t)+\frac{j}{\pi t}\right)*u_{k}(t)$ represents the analytic signal of $u_{k}(t)$. The objective function in Equation (3) takes smaller values when the bandwidth around the center angular frequency, $\omega_{k}$, of each mode, $u_{k}(t)$, in the frequency distribution is small. In the actual algorithm, the following augmented Lagrangian is defined to loosen the constrained variational optimization problem (3). That is, the solution to Equation (3) is sought as the saddle point of the following augmented Lagrangian (4):
@@
This problem (4) can be solved using the alternate direction method of multipliers (ADMM) Boyd2011 . $K$ and $\alpha$ are parameters set by the user, allowing the adjustment of the number of modes and the penalty factor, respectively.
@@ 2.2 Multivariate Variational Mode Decomposition
MVMD is an extension of VMD that applies to multiple variables. It expresses multivariate data as a sum of $K$ time-series patterns. Here, the input data are denoted as $y_{c}(t)$, where $c=1,2,\cdots,C$, and subscript $c$ indicates the $c$-th variable. The $k$-th MVMD mode corresponding to $y_{c}(t)$ is expressed as $u_{k,c}(t)$. In MVMD, the MVMD mode, $u_{k,c}(t)$, and the center angular frequency, $\omega_{k}$, are computed as solutions to the following optimization problem:
@@ 2.3 Alternate Direction Method of Multipliers for Multivariate Variational Mode Decomposition
         Update center frequency $\hat{\omega}_{k}$:
@@
In this study, the saddle point of (6) was solved using the ADMM method, referring to Ref. Rehman2019 . ADMM involves solving sub-optimization problems and iteratively updating the modes, center angular frequencies, and Lagrange multipliers. The solutions to the sub-optimization problems are given below.
@@ 2.4.3 Intrinsic coherent structure
That is, by applying MVMD, this method extracts the fluctuation components near the center frequencies of each $k$ from the POD expansion coefficients. Then, the spatiotemporal patterns represented by the $k$-th MVMD mode and the POD basis, namely the nonstationary (or stationary) coherent structure extracted using the VMD-NCS analysis, are computed by the following equation:
@@
Equations (17)–(18) explicitly state that the ICS ($\psi_{k}(\bm{x},t)$) is a function of position $\bm{x}$ and time $t$. Thus, in contrast to the case in the conventional modal decomposition (1), the spatial distribution of the ICS can vary with time.
@@ 2.4.4 Amplitude of the intrinsic coherent structure
The amplitude of the $k$-th ICS, $\|\psi_{k}(t)\|$, can be easily computed due to the orthogonality of the POD basis, $P$, as follows:
$\|\psi_{k}(t)\|=\sqrt{\sum_{i=1}^{r}u_{k,i}(t)^{2}}$ (19)
@@ 2.4.5 Setting parameters
In MVMD, there are two user-setting parameters: $\alpha$, which regulates the penalty term in Equation (6), and $K$, which represents the number of ICSs. A large $\alpha$ results in a narrow bandwidth effect. Accordingly, the extracted ICSs capture phenomena with relatively narrow frequencies, i.e., high periodicity, while failing to represent nonperiodic variations. Conversely, reducing parameter $\alpha$ weakens the periodicity, and if $\alpha$ becomes overly small, the temporal coherence of ICSs may be lost. Notably, a high $K$ and a low $\alpha$ increase the difficulty of comprehending the phenomena. Conversely, if the $K$ is overly low or $\alpha$ is overly high, it might be unfeasible to represent the input data through the superposition of ICSs, increasing the reconstruction error (10). Therefore, it is considered important to set parameters $K$ and $\alpha$ within ranges that maintain the ease with which the phenomena are understood and where a sufficiently small reconstruction error can be achieved. The parameter dependency when applying the VMD-NCS analysis to actual flow fields is discussed in Sections 3 and 4.
@@ 3.1 Problem Setting and Input data
As the first validation case, the transient growth phenomena around a two-dimensional square cylinder were analyzed. The input data consisted of temporal distributions of $x$- and $y$-direction velocities obtained through a two-dimensional numerical fluid simulation. In the numerical fluid simulation, first, we calculated, using the selective frequency damping method Akervik2006 , an unstable steady-state solution, in which twin vortices appeared behind the cylinder. Subsequently, by evolving this steady-state solution from time $t=0$ as the initial condition, temporal data capturing the transient growth phenomena from the unstable steady-state to the periodic shedding of Kármán vortices were acquired. The freestream Reynolds number based on the length of one side of the square cylinder was 120, and the freestream Mach number was 0.2. For details on this numerical fluid simulation, please refer to Ref. Ohmichi2021 .
@@ 3.2.1 Extracted coherent structures ($K=5,~{}\alpha=100$)
First, the ICSs at $K=5$ and $\alpha=100$ are evaluated. Fig. 2 shows the temporal evolution of the ICS amplitudes, $\|\bm{\phi}_{k}(t)\|$, and the time-frequency distribution obtained by the Hilbert transform of the ICSs with respect to the $x$-direction velocity at position $(x,~{}y)=(5,~{}0.5)$. Setting $K=5$ yields five ICSs. Among these obtained ICSs, three exhibited significant amplitudes at $St_{c}=0.003$, 0.112, and 0.147 (Fig. 2(a)). From Fig. 2(a), it is evident that the ICS of $St_{c}=0.112$ became dominant during the transient growth period of the flow (i.e., it had a large amplitude), whereas the ICS of $St_{c}=0.147$ became dominant during the subsequent periodic shedding of the Kármán vortices. The spatial distribution and amplitude variation of these ICSs with time are depicted in Fig. 3. The figure illustrates the growth of the ICS of $St_{c}=0.112$, representing transient growth, followed by the appearance of the ICS of $St_{c}=0.147$, representing the periodic shedding of the Kármán vortices. The figure also shows the subsequent disappearance of the ICS of $St_{c}=0.112$. Moreover, for the ICS of $St_{c}=0.112$, the vortex shedding initially manifested at the far-wake region of the cylinder ($t=112.5$) and gradually approached the near-wake region of the cylinder. Note that such spatial-distribution changes cannot be represented as a single mode using conventional methods in the form of Equation (1). This result shows that these spatial-distribution changes can be represented as a single ICS using the VMD-NCS analysis method.
Furthermore, on examining the results of the time-frequency analysis (Fig. 2(b)) it was confirmed that, as intended, each ICS exhibited time-varying frequency appearing around their center frequencies, and the frequency and amplitude gradually varied over time.
@@
Figure 5: The spatial distribution of the $x$-direction velocity component for the (a) first, (b) third, (c) fifth, (d) seventh, and (e) ninth POD modes (left); temporal evolution of the expansion coefficients for each POD mode (middle); MVMD modes obtained by decomposing the expansion coefficients (right).
Fig. 5 illustrates the spatial distributions of the POD modes, $\bm{\phi}_{i}$, and the temporal evolution of the corresponding expansion coefficients, $\tilde{q}_{i}(t)$. Additionally, the right column presents the MVMD modes, $u_{k,i}(t)$, obtained by decomposing these POD coefficients (16). On observing the temporal evolution of the POD expansion coefficients, it is evident that each POD mode comprises components of various phenomena, including transient changes, periodic variations, and gradual temporal changes in the background flow. For instance, the coefficients of the first POD mode contained a mixture of slowly varying components and periodic fluctuations. Similarly, coefficients of other modes exhibited variations representing the transient growth of the flow field and periodic fluctuations within a single mode. Conversely, the MVMD modes effectively decomposed these into temporal variations, corresponding to individual phenomena. In the VMD-NCS analysis, the MVMD method enables the extraction of individual phenomena from the temporal evolution of the POD mode expansion coefficients, and superimposing the POD modes for each extracted phenomena enables the extraction of spatiotemporal patterns of nonstationary phenomena as ICSs.

### v2

@@ Abstract
The conventional modal analysis techniques face difficulties in handling nonstationary phenomena, such as transient, nonperiodic, or intermittent phenomena. This paper presents a variational mode decomposition–based nonstationary coherent structure (VMD-NCS) analysis that enables the extraction and analysis of coherent structures in the case of nonstationary phenomena from high-dimensional spatiotemporal data. The VMD-NCS analysis decomposes the input spatiotemporal data into intrinsic coherent structures (ICSs) that represent nonstationary spatiotemporal patterns and exhibit coherence in both spatial and temporal directions. Furthermore, unlike many conventional modal analysis techniques, the proposed method accounts for the temporal changes in the spatial distribution with time. The performance of the VMD-NCS analysis was validated based on the transient growth phenomena in the flow around a cylinder. It was confirmed that the temporal changes in the spatial distribution, depicting the transient growth of vortex shedding where fluctuations arising in the far-wake region gradually approach the near-wake region, were represented as a single ICS. Furthermore, in the analysis of the quasi-periodic flow field around a pitching airfoil, the temporal changes in the spatial distribution and the amplitude of vortex shedding behind the airfoil, influenced by the pitching motion of the airfoil, were captured as a single ICS. Additionally, the impact of two parameters that control the number of ICSs ($K$) and the penalty factor related to the temporal coherence ($\alpha$), was investigated. The results revealed that $K$ has a significant impact on the VMD-NCS analysis results. In the case of a relatively high $K$, the VMD-NCS analysis tends to extract more periodic spatiotemporal patterns resembling the results of dynamic mode decomposition. In the case of a small $K$, it tends to extract more nonstationary spatiotemporal patterns.
@@
journal: Aerospace Science and Technology
@@ 1 Introduction
Generally, many physical phenomena generate complex spatiotemporal structures owing to their nonlinearity. In recent years, techniques for extracting spatiotemporal patterns from high-dimensional time-series data obtained through computational fluid dynamics or experiments have been actively researched in the fluid dynamics field. These techniques aim to elucidate the physically significant fluid structures and their temporal behaviors within complex flow fields. Additionally, they are employed to model the flow field for prediction and control [1, 2, 3, 4]. The most common analytical methods are proper orthogonal decomposition (POD) [5, 6] and dynamic mode decomposition (DMD) [7]. POD identifies the dominant spatiotemporal patterns by computing orthogonal bases that minimize the reconstruction error when reconstructing the original data. DMD represents the flow field’s temporal evolution from one time to the next using a linear operator, thereby extracting modes as eigenvalues and eigenvectors of the operator. Due to the linear-operator representation, each mode exhibits a linear growth in periodic oscillation amplitudes, simplifying the physical interpretation of the obtained modes. These techniques have found extensive applications in the extraction of spatiotemporal patterns from complex flow fields, such as jets [8, 9, 10], wakes [11, 12, 13], and flows involving shockwaves [14, 15, 16]. Moreover, substantial research efforts have been devoted to developing advanced methodologies, encompassing algorithms tailored for large and streaming datasets [17, 18, 19], noise-robust algorithms [20, 21, 22, 23, 24], data-reconstruction methods from noisy [25, 26, 27] or incomplete [28, 29] data, and stability analysis methods [30, 31].
@@
However, existing modal analysis techniques encounter difficulties in handling transient, nonperiodic, or intermittent phenomena (nonstationary phenomena). For instance, the DMD method represents each mode’s temporal behavior as a linear growth and periodic oscillation. Thus, it is unable to capture temporal variations in amplitude and frequency inherent in nonstationary phenomena. The POD method is frequently limited by the coexistence of various phenomena within a single mode, complicating its physical interpretation. Additionally, most existing modal-decomposition methods rely on input data representations of the following form [32]:
$f(\bm{x},t)=\sum_{i=1}^{N}{w_{i}(t)\phi_{i}(\bm{x})}+\epsilon(\bm{x},t)$ (1)
@@
Here, the input data are represented as a superposition of $N$ space-dependent modes, $\phi_{i}$. The coefficients, $w_{i}$, are expansion coefficients that vary with time. $\epsilon$ denotes the expansion residual. Equation (1) indicates that the spatial distribution of mode $\phi_{i}$ remains unchanged with time. However, many nonstationary phenomena, such as the transient growth of a cylinder wake, involve spatial structures that vary with time. Therefore, to extract such phenomena as a single mode, a modal expansion using ‘flexible’ modes capable of changing spatial structures is required, as noted by Noack [32].
@@
In the field of signal processing, variational mode decomposition (VMD) [33] and its multivariate counterpart, multivariate VMD (MVMD) [34], have recently found applications as analytical tools for addressing nonstationary time-varying data. VMD tackles optimization problems by incorporating penalty terms related to the frequency bandwidth of modes, thereby extracting the coherent fluctuation patterns (modes) within the input signal. Unlike DMD, VMD does not assume single periodic variations and can represent nonlinear changes in amplitude. Each mode exhibits a distinct frequency spectrum with different center frequencies. Additionally, by extracting modes with narrow frequency bandwidths, it becomes feasible to isolate similar frequency variations as one coherent phenomenon, enhancing the ease of deriving physical interpretations. However, directly applying VMD algorithms to high-dimensional time-series data, such as fluid data, is challenging.
@@
Only a few studies have attempted to develop modal analysis techniques for nonstationary fluid phenomena. The study by Liao et al. [35] is the sole study focusing on an approach based on VMD. They proposed reduced-order VMD (RVMD) inspired by the Hilbert–Huang transform and VMD. However, RVMD is a modal-decomposition method that is represented as Equation (1), which cannot capture the temporal evolution of spatial distributions of nonstationary phenomena as a single mode. In the approach based on DMD, called multi-resolution DMD (mrDMD) [36], different dynamic modes are extracted for each temporal domain of data based on the repetitive removal of low-frequency components and data segmentation. The frequency-time analysis using the spectral POD is also based on data segmentation [37]. The conditional space–time POD method [38] extracts dominant spatiotemporal patterns from a set of event realizations. The modes of the conditional space-time POD exhibit coherent structures in both spatial and temporal directions. Recently, Stahl et al. [31] reported that convective instability can be analyzed by applying DMD to conditional space-time POD modes.
@@ 2.1 Variational Mode Decomposition
where $K$ is the number of modes. VMD is a data-driven modal-decomposition method, and each mode, $u_{k}(t)$, is computed from the input signal, $y(t)$. The purpose of VMD is to represent the input signal as a sum of modes $u_{k}(t)$, each with narrow-band frequency components. To achieve this, the modes are determined such that the sum of the bandwidths of all modes is minimized. To estimate the bandwidth of each mode, the following approach was proposed in the orignial VMD algorithm: 1) Compute the analytic signal for each mode via the Hilbert transform, resulting in a single-sided frequency spectrum. 2) Shift the frequency spectrum of each mode to the baseband by multiplying it with a complex exponential at the respective estimated center frequency. 3) Estimate the bandwidth by taking the squared L2 norm of the gradient of the harmonically mixed signal obtained in the previous step. This formulation leads to the following constrained variational problem, aiming to minimize the bandwidth of the frequency distribution of each VMD mode and the reconstruction error:
@@
Here, $\left(\delta(t)+\frac{j}{\pi t}\right)*u_{k}(t)$ represents the analytic signal of $u_{k}(t)$. The $\delta(\cdot)$ and * denote the Dirac delta function and the convolution, respectively. The objective function in Equation (3) takes smaller values when the bandwidth around the center angular frequency, $\omega_{k}$, of each mode, $u_{k}(t)$, in the frequency distribution is small. Please refer to Ref. [33] for further mathematical details for the estimation of the bandwidth in Equation (3).
In the actual algorithm, the following augmented Lagrangian is defined to loosen the constrained variational optimization problem (3). That is, the solution to Equation (3) is sought as the saddle point of the following augmented Lagrangian (4):
@@
This problem (4) can be solved using the alternate direction method of multipliers (ADMM) [39]. $K$ and $\alpha$ are parameters set by the user, allowing the adjustment of the number of modes and the penalty factor, respectively.
@@ 2.2 Multivariate Variational Mode Decomposition
MVMD is an extension of VMD that applies to multiple variables. It expresses multivariate data as a sum of $K$ time-series patterns. Here, the input data are denoted as $y_{c}(t)$, where $c=1,2,\cdots,C$, and subscript $c$ indicates the $c$-th variable. The $k$-th MVMD mode corresponding to the $c$-th variable $y_{c}(t)$ is expressed as $u_{k,c}(t)$. In MVMD, the MVMD mode, $u_{k,c}(t)$, and the center angular frequency, $\omega_{k}$, are computed as solutions to the following optimization problem:
@@ 2.3 Alternate Direction Method of Multipliers for Multivariate Variational Mode Decomposition
         Update center angular frequency $\hat{\omega}_{k}$:
@@
In this study, the saddle point of (6) was solved using the ADMM method, referring to Ref. [34]. ADMM involves solving sub-optimization problems and iteratively updating the modes, center angular frequencies, and Lagrange multipliers. The solutions to the sub-optimization problems are given below.
@@ 2.4.3 Intrinsic coherent structure
That is, by applying MVMD, this method decomposes the POD expansion coefficients into separate temporal patterns, each associated with a center frequency corresponding to $k$. Then, the spatiotemporal patterns represented by the $k$-th MVMD mode and the POD basis, namely the nonstationary (or stationary) coherent structure extracted using the VMD-NCS analysis, are computed by the following equation:
@@
Note that, Equations (17)–(18) explicitly state that the ICS ($\psi_{k}(\bm{x},t)$) is a function of position $\bm{x}$ and time $t$. Thus, in contrast to the case in the conventional modal decomposition (1), the spatial distribution of the ICS can vary with time.
@@ 2.4.4 Amplitude of the intrinsic coherent structure
The amplitude of the $k$-th ICS at time $t=t_{j}$, $\|\psi_{k}(t_{j})\|$, can be easily computed due to the orthogonality of the POD basis, $P$, as follows:
$\|\psi_{k}(t_{j})\|=\sqrt{\sum_{i=1}^{r}u_{k,i}(t_{j})^{2}}$ (19)
@@ 2.4.5 Setting parameters
In MVMD, there are two user-setting parameters: $\alpha$, which regulates the penalty term in Equation (6), and $K$, which represents the number of ICSs. A large $\alpha$ results in a narrow bandwidth effect. Accordingly, the extracted ICSs capture phenomena with relatively narrow frequencies, i.e., high periodicity, while failing to represent nonperiodic variations. Conversely, reducing parameter $\alpha$ weakens the periodicity. If $\alpha$ becomes overly small, the temporal coherence of ICSs may be lost. Notably, a high $K$ and a low $\alpha$ increase the difficulty of comprehending the phenomena. Conversely, if the $K$ is overly low or $\alpha$ is overly high, it might be unfeasible to represent the input data through the superposition of ICSs, increasing the reconstruction error (10). Therefore, it is important to set parameters $K$ and $\alpha$ within a range that ensures the extracted ICSs are easily interpretable while maintaining a sufficiently small reconstruction error. The parameter dependency when applying the VMD-NCS analysis to actual flow fields is discussed in Sections 3 and 4.
@@ 3.1 Problem Setting and Input data
As the first validation case, the transient growth phenomena around a two-dimensional square cylinder were analyzed. The input data consisted of temporal distributions of $x$- and $y$-direction velocities obtained through a two-dimensional numerical fluid simulation. In the numerical fluid simulation, we first employed the selective frequency damping method [40] to compute an unstable steady-state solution, which exhibited twin vortices behind the cylinder. Subsequently, by evolving this steady-state solution from time $t=0$ as the initial condition, temporal data capturing the transient growth phenomena from the unstable steady-state to the periodic shedding of Kármán vortices were acquired. The freestream Reynolds number based on the length of one side of the square cylinder was 120, and the freestream Mach number was 0.2. For details on this numerical fluid simulation, please refer to Ref. [41].
@@ 3.2.1 Extracted coherent structures ($K=5,~{}\alpha=100$)
First, the ICSs at $K=5$ and $\alpha=100$ are evaluated. Fig. 2 shows the temporal evolution of the ICS amplitudes, $\|\bm{\psi}_{k}(t)\|$, and the time-frequency distribution obtained by the Hilbert transform of the ICSs with respect to the $x$-direction velocity at position $(x,~{}y)=(5,~{}0.5)$. Setting $K=5$ yields five ICSs. Among these obtained ICSs, three exhibited significant amplitudes at $St_{c}=0.003$, 0.112, and 0.147 (Fig. 2(a)). Here, $St_{c}$ denotes the nondimensionalized center frequency of each ICS. From Fig. 2(a), it is evident that the ICS of $St_{c}=0.112$ became dominant during the transient growth period of the flow (i.e., it had a large amplitude), whereas the ICS of $St_{c}=0.147$ became dominant during the subsequent periodic shedding of the Kármán vortices. The spatial distribution and amplitude variation of these ICSs with time are depicted in Fig. 3. The figure illustrates the growth of the ICS of $St_{c}=0.112$, representing transient growth, followed by the appearance of the ICS of $St_{c}=0.147$, representing the periodic shedding of the Kármán vortices. The figure also shows the subsequent disappearance of the ICS of $St_{c}=0.112$. Moreover, for the ICS of $St_{c}=0.112$, the vortex shedding initially manifested at the far-wake region of the cylinder ($t=112.5$) and gradually approached the near-wake region of the cylinder. Note that such spatial-distribution changes cannot be represented as a single mode using conventional methods, including POD and DMD, which are in the form of Equation (1). This result demonstrates that these spatial-distribution changes can be represented as a single ICS using the VMD-NCS analysis method, highlighting its advantage over conventional methods in capturing the time-varying nature of nonstationary phenomena.
Furthermore, on examining the results of the time-frequency analysis (Fig. 2(b)), it was confirmed that each ICS exhibited time-varying frequency appearing around their center frequencies, and the frequency and amplitude gradually varied over time.
@@
Figure 5: The spatial distribution of the $x$-direction velocity component for the (a) first, (b) third, (c) fifth, (d) seventh, and (e) ninth POD modes.
Figure 6: The temporal evolution of the expansion coefficients (left) and MVMD modes obtained by decomposing the expansion coefficients (right) for the (a) first, (b) third, (c) fifth, (d) seventh, and (e) ninth POD modes.
Fig. 5 shows the spatial distributions of the POD modes $\bm{\phi}_{i}$. Additionally, Fig. 6 presents the temporal evolution of the POD coefficients $\tilde{q}_{i}(t)$, and the MVMD modes $u_{k,i}(t)$ obtained by decomposing these POD coefficients (16). The spatial distribution of the POD modes appears to resemble the distribution of ICSs at some time instances shown in Figs. 3 and 4. However, on observing the temporal evolution of the POD expansion coefficients, it is evident that each POD mode comprises components of various phenomena, including transient changes, periodic variations, and gradual temporal changes in the background flow. For instance, the coefficients of the first POD mode contained a mixture of slowly varying components and periodic fluctuations. Similarly, coefficients of other modes exhibited variations representing the transient growth of the flow field and periodic fluctuations within a single mode. Conversely, the MVMD modes effectively decomposed these into temporal variations, corresponding to individual phenomena. In the VMD-NCS analysis, the MVMD method enables the extraction of individual phenomena from the temporal evolution of the POD mode expansion coefficients, and superimposing the POD modes for each extracted phenomena enables the extraction of spatiotemporal patterns of nonstationary phenomena as ICSs.